Required information (The following information applies to the questions displayed below.] Sweeten Company had no jobs in pro2.  How much manufacturing overhead was applied to Job P and how much was applied to Job Q?
Answer #1

Solution 2:

Total estimated Overhead = (2500*1.40) + (1500*2.20) + $25000 = $31,800

Predetermined Overhead rate = $31800/ 4000 = $7.95 per Machine hour

Overhead assigned to Job P = 2300* $7.95 = $18,285

Overhead assigned to Job Q = 1700* $7.95 = $13,515